\begin{thebibliography}{99}





\bibitem{inflation1}
 A.~D.~Linde,
 %``Inflationary Cosmology,''
 Lect.\ Notes Phys.\ {\bf 738} (2008) 1
 %%%doi:10.1007/978-3-540-74353-8_1
 [arXiv:0705.0164 [hep-th]].
 %%CITATION = %%doi:10.1007/978-3-540-74353-8_1;%%
 %336 citations counted in INSPIRE as of 24 Dec 2016

\bibitem{inflation2} D.~S.~Gorbunov and V.~A.~Rubakov,
``Introduction to the theory of the early universe: Cosmological
perturbations and inflationary theory,'' Hackensack, USA: World
Scientific (2011) 489 p;
%6 citations counted in INSPIRE as of 14 Feb 2015
%
%V. Mukhanov, Physical foundations of cosmology, Cambridge, UK: Univ. Pr. (2005) 421 p; D. S. Gorbunov, V. A. Rubakov, Introduction to the theory of the early Universe: Cosmological perturbations and inflationary theory, Hackensack, USA, World Scientific (2011) 489 p


\bibitem{inflation3}A.~Linde,
%``Inflationary Cosmology after Planck 2013,''
arXiv:1402.0526 [hep-th];
%\cite{Lyth:1998xn}
%\bibitem{Lyth:1998xn}


\bibitem{inflation4}D.~H.~Lyth and A.~Riotto,
%``Particle physics models of inflation and the cosmological density perturbation,''
Phys.\ Rept.\  {\bf 314} (1999) 1 [hep-ph/9807278].



%\cite{Odintsov:2023weg}
\bibitem{inflation5}
S.~D.~Odintsov, V.~K.~Oikonomou, I.~Giannakoudi, F.~P.~Fronimos
and E.~C.~Lymperiadou,
%``Recent Advances in Inflation,''
Symmetry \textbf{15} (2023) no.9, 1701 doi:10.3390/sym15091701
[arXiv:2307.16308 [gr-qc]].
%18 citations counted in INSPIRE as of 18 Nov 2023



%\cite{CMB-S4:2016ple}
\bibitem{CMB-S4:2016ple}
K.~N.~Abazajian \textit{et al.} [CMB-S4],
%``CMB-S4 Science Book, First Edition,''
[arXiv:1610.02743 [astro-ph.CO]].
%1048 citations counted in INSPIRE as of 26 Oct 2021



%\cite{SimonsObservatory:2019qwx}
\bibitem{SimonsObservatory:2019qwx}
M.~H.~Abitbol \textit{et al.} [Simons Observatory],
%``The Simons Observatory: Astro2020 Decadal Project Whitepaper,''
Bull. Am. Astron. Soc. \textbf{51} (2019), 147 [arXiv:1907.08284
[astro-ph.IM]].
%42 citations counted in INSPIRE as of 26 Oct 2021



%\cite{Hild:2010id}
\bibitem{Hild:2010id}
S.~Hild, M.~Abernathy, F.~Acernese, P.~Amaro-Seoane, N.~Andersson,
K.~Arun, F.~Barone, B.~Barr, M.~Barsuglia and M.~Beker, \textit{et
al.}
%``Sensitivity Studies for Third-Generation Gravitational Wave Observatories,''
Class. Quant. Grav. \textbf{28} (2011), 094013
doi:10.1088/0264-9381/28/9/094013 [arXiv:1012.0908 [gr-qc]].
%353 citations counted in INSPIRE as of 28 Apr 2021




%\cite{Baker:2019nia}
\bibitem{Baker:2019nia}
J.~Baker, J.~Bellovary, P.~L.~Bender, E.~Berti, R.~Caldwell,
J.~Camp, J.~W.~Conklin, N.~Cornish, C.~Cutler and R.~DeRosa,
\textit{et al.}
%``The Laser Interferometer Space Antenna: Unveiling the Millihertz Gravitational Wave Sky,''
[arXiv:1907.06482 [astro-ph.IM]].
%39 citations counted in INSPIRE as of 28 Apr 2021


%\cite{Smith:2019wny}
\bibitem{Smith:2019wny}
T.~L.~Smith and R.~Caldwell,
%``LISA for Cosmologists: Calculating the Signal-to-Noise Ratio for Stochastic and Deterministic Sources,''
Phys. Rev. D \textbf{100} (2019) no.10, 104055
doi:10.1103/PhysRevD.100.104055 [arXiv:1908.00546 [astro-ph.CO]].
%24 citations counted in INSPIRE as of 28 Apr 2021


%\cite{Crowder:2005nr}
\bibitem{Crowder:2005nr}
J.~Crowder and N.~J.~Cornish,
%``Beyond LISA: Exploring future gravitational wave missions,''
Phys. Rev. D \textbf{72} (2005), 083005
doi:10.1103/PhysRevD.72.083005 [arXiv:gr-qc/0506015 [gr-qc]].
%242 citations counted in INSPIRE as of 28 Apr 2021


%\cite{Smith:2016jqs}
\bibitem{Smith:2016jqs}
T.~L.~Smith and R.~Caldwell,
%``Sensitivity to a Frequency-Dependent Circular Polarization in an Isotropic Stochastic Gravitational Wave Background,''
Phys. Rev. D \textbf{95} (2017) no.4, 044036
doi:10.1103/PhysRevD.95.044036 [arXiv:1609.05901 [gr-qc]].
%36 citations counted in INSPIRE as of 28 Apr 2021



%\cite{Seto:2001qf}
\bibitem{Seto:2001qf}
N.~Seto, S.~Kawamura and T.~Nakamura,
%``Possibility of direct measurement of the acceleration of the universe using 0.1-Hz band laser interferometer gravitational wave antenna in space,''
Phys. Rev. Lett. \textbf{87} (2001), 221103
doi:10.1103/PhysRevLett.87.221103 [arXiv:astro-ph/0108011
[astro-ph]].
%468 citations counted in INSPIRE as of 28 Apr 2021


%\cite{Kawamura:2020pcg}
\bibitem{Kawamura:2020pcg}
S.~Kawamura, M.~Ando, N.~Seto, S.~Sato, M.~Musha, I.~Kawano,
J.~Yokoyama, T.~Tanaka, K.~Ioka and T.~Akutsu, \textit{et al.}
%``Current status of space gravitational wave antenna DECIGO and B-DECIGO,''
[arXiv:2006.13545 [gr-qc]].
%26 citations counted in INSPIRE as of 28 Apr 2021



%\cite{Bull:2018lat}
\bibitem{Bull:2018lat}
A.~Weltman, P.~Bull, S.~Camera, K.~Kelley, H.~Padmanabhan,
J.~Pritchard, A.~Raccanelli, S.~Riemer-S\o{}rensen, L.~Shao and
S.~Andrianomena, \textit{et al.}
%``Fundamental physics with the Square Kilometre Array,''
Publ. Astron. Soc. Austral. \textbf{37} (2020), e002
doi:10.1017/pasa.2019.42 [arXiv:1810.02680 [astro-ph.CO]].
%82 citations counted in INSPIRE as of 28 Apr 2021



%\cite{LISACosmologyWorkingGroup:2022jok}
\bibitem{LISACosmologyWorkingGroup:2022jok}
P.~Auclair \textit{et al.} [LISA Cosmology Working Group],
%``Cosmology with the Laser Interferometer Space Antenna,''
Living Rev. Rel. \textbf{26} (2023) no.1, 5
doi:10.1007/s41114-023-00045-2 [arXiv:2204.05434 [astro-ph.CO]].
%183 citations counted in INSPIRE as of 18 Nov 2023


%\cite{Vagnozzi:2022qmc}
\bibitem{Vagnozzi:2022qmc}
S.~Vagnozzi and A.~Loeb,
%``The Challenge of Ruling Out Inflation via the Primordial Graviton Background,''
Astrophys. J. Lett. \textbf{939} (2022) no.2, L22
doi:10.3847/2041-8213/ac9b0e [arXiv:2208.14088 [astro-ph.CO]].
%24 citations counted in INSPIRE as of 20 Jan 2024




\bibitem{nanograv}

%\cite{NANOGrav:2023gor}
%\bibitem{NANOGrav:2023gor}
G.~Agazie \textit{et al.} [NANOGrav],
%``The NANOGrav 15 yr Data Set: Evidence for a Gravitational-wave Background,''
Astrophys. J. Lett. \textbf{951} (2023) no.1, L8
doi:10.3847/2041-8213/acdac6 [arXiv:2306.16213 [astro-ph.HE]].
%318 citations counted in INSPIRE as of 18 Nov 2023


%EPTA

%\cite{Antoniadis:2023ott}
\bibitem{Antoniadis:2023ott}
J.~Antoniadis, P.~Arumugam, S.~Arumugam, S.~Babak, M.~Bagchi,
A.~S.~B.~Nielsen, C.~G.~Bassa, A.~Bathula, A.~Berthereau and
M.~Bonetti, \textit{et al.}
%``The second data release from the European Pulsar Timing Array III. Search for gravitational wave signals,''
[arXiv:2306.16214 [astro-ph.HE]].
%0 citations counted in INSPIRE as of 29 Jun 2023



%\cite{Reardon:2023gzh}
\bibitem{Reardon:2023gzh}
D.~J.~Reardon, A.~Zic, R.~M.~Shannon, G.~B.~Hobbs, M.~Bailes,
V.~Di Marco, A.~Kapur, A.~F.~Rogers, E.~Thrane and J.~Askew,
\textit{et al.}
%``Search for an isotropic gravitational-wave background with the Parkes Pulsar Timing Array,''
doi:10.3847/2041-8213/acdd02 [arXiv:2306.16215 [astro-ph.HE]].
%0 citations counted in INSPIRE as of 29 Jun 2023


%\cite{Xu:2023wog}
\bibitem{Xu:2023wog}
H.~Xu, S.~Chen, Y.~Guo, J.~Jiang, B.~Wang, J.~Xu, Z.~Xue,
R.~N.~Caballero, J.~Yuan and Y.~Xu, \textit{et al.}
%``Searching for the nano-Hertz stochastic gravitational wave background with the Chinese Pulsar Timing Array Data Release I,''
doi:10.1088/1674-4527/acdfa5 [arXiv:2306.16216 [astro-ph.HE]].
%0 citations counted in INSPIRE as of 29 Jun 2023


%\cite{NANOGrav:2023gor}
\bibitem{NANOGrav:2023gor}
G.~Agazie \textit{et al.} [NANOGrav],
%``The NANOGrav 15-year Data Set: Evidence for a Gravitational-Wave Background,''
 [arXiv:2306.16221 [astro-ph.HE]].
%5 citations counted in INSPIRE as of 29 Jun 2023



%\cite{Sampson:2015ada}
\bibitem{Sampson:2015ada}
L.~Sampson, N.~J.~Cornish and S.~T.~McWilliams,
%``Constraining the Solution to the Last Parsec Problem with Pulsar Timing,''
Phys. Rev. D \textbf{91} (2015) no.8, 084055
doi:10.1103/PhysRevD.91.084055 [arXiv:1503.02662 [gr-qc]].
%48 citations counted in INSPIRE as of 22 Jun 2023



\bibitem{sunnynew}
S.~Vagnozzi,
%``Inflationary interpretation of the stochastic gravitational wave background signal detected by pulsar timing array experiments,''
JHEAp \textbf{39} (2023), 81-98 doi:10.1016/j.jheap.2023.07.001
[arXiv:2306.16912 [astro-ph.CO]].
%59 citations counted in INSPIRE as of 06 Sep 2023




%\cite{Oikonomou:2023qfz}
\bibitem{Oikonomou:2023qfz}
V.~K.~Oikonomou,
%``Flat energy spectrum of primordial gravitational waves versus peaks and the NANOGrav 2023 observation,''
Phys. Rev. D \textbf{108} (2023) no.4, 043516
doi:10.1103/PhysRevD.108.043516 [arXiv:2306.17351 [astro-ph.CO]].
%14 citations counted in INSPIRE as of 29 Aug 2023



%\cite{Cai:2023dls}
\bibitem{Cai:2023dls}
Y.~F.~Cai, X.~C.~He, X.~Ma, S.~F.~Yan and G.~W.~Yuan,
%``Limits on scalar-induced gravitational waves from the stochastic background by pulsar timing array observations,''
[arXiv:2306.17822 [gr-qc]].
%0 citations counted in INSPIRE as of 04 Jul 2023


%\cite{Han:2023olf}
\bibitem{Han:2023olf}
C.~Han, K.~P.~Xie, J.~M.~Yang and M.~Zhang,
%``Self-interacting dark matter implied by nano-Hertz gravitational waves,''
[arXiv:2306.16966 [hep-ph]].
%3 citations counted in INSPIRE as of 04 Jul 2023



%\cite{Guo:2023hyp}
\bibitem{Guo:2023hyp}
S.~Y.~Guo, M.~Khlopov, X.~Liu, L.~Wu, Y.~Wu and B.~Zhu,
%``Footprints of Axion-Like Particle in Pulsar Timing Array Data and JWST Observations,''
[arXiv:2306.17022 [hep-ph]].
%3 citations counted in INSPIRE as of 04 Jul 2023


%\cite{Yang:2023aak}
\bibitem{Yang:2023aak}
J.~Yang, N.~Xie and F.~P.~Huang,
%``Nano-Hertz stochastic gravitational wave background as hints of ultralight axion particles,''
[arXiv:2306.17113 [hep-ph]].
%3 citations counted in INSPIRE as of 04 Jul 2023


%\cite{Addazi:2023jvg}
\bibitem{Addazi:2023jvg}
A.~Addazi, Y.~F.~Cai, A.~Marciano and L.~Visinelli,
%``Have pulsar timing array methods detected a cosmological phase transition?,''
[arXiv:2306.17205 [astro-ph.CO]].
%0 citations counted in INSPIRE as of 04 Jul 2023


%\cite{Li:2023bxy}
\bibitem{Li:2023bxy}
S.~P.~Li and K.~P.~Xie,
%``Collider test of nano-Hertz gravitational waves from pulsar timing arrays,''
Phys. Rev. D \textbf{108} (2023) no.5, 055018
doi:10.1103/PhysRevD.108.055018 [arXiv:2307.01086 [hep-ph]].
%25 citations counted in INSPIRE as of 20 Sep 2023


%\cite{Niu:2023bsr}
\bibitem{Niu:2023bsr}
X.~Niu and M.~H.~Rahat,
%``NANOGrav signal from axion inflation,''
[arXiv:2307.01192 [hep-ph]].
%0 citations counted in INSPIRE as of 04 Jul 2023



%\cite{Yang:2023qlf}
\bibitem{Yang:2023qlf}
A.~Yang, J.~Ma, S.~Jiang and F.~P.~Huang,
%``Implication of nano-Hertz stochastic gravitational wave on dynamical dark matter through a first-order phase transition,''
[arXiv:2306.17827 [hep-ph]].
%10 citations counted in INSPIRE as of 13 Jul 2023




%\cite{Datta:2023vbs}
\bibitem{Datta:2023vbs}
S.~Datta,
%``Inflationary gravitational waves, pulsar timing data and low-scale-leptogenesis,''
[arXiv:2307.00646 [hep-ph]].
%5 citations counted in INSPIRE as of 13 Jul 2023

%\cite{Du:2023qvj}
\bibitem{Du:2023qvj}
X.~K.~Du, M.~X.~Huang, F.~Wang and Y.~K.~Zhang,
%``Did the nHZ Gravitational Waves Signatures Observed By NANOGrav Indicate Multiple Sector SUSY Breaking?,''
[arXiv:2307.02938 [hep-ph]].
%5 citations counted in INSPIRE as of 14 Jul 2023





%\cite{Salvio:2023ynn}
\bibitem{Salvio:2023ynn}
A.~Salvio,
%``Supercooling in Radiative Symmetry Breaking: Theory Extensions, Gravitational Wave Detection and Primordial Black Holes,''
[arXiv:2307.04694 [hep-ph]].
%1 citations counted in INSPIRE as of 14 Jul 2023

%\cite{Yi:2023mbm}
\bibitem{Yi:2023mbm}
Z.~Yi, Q.~Gao, Y.~Gong, Y.~Wang and F.~Zhang,
%``The waveform of the scalar induced gravitational waves in light of Pulsar Timing Array data,''
[arXiv:2307.02467 [gr-qc]].
%7 citations counted in INSPIRE as of 14 Jul 2023


%\cite{You:2023rmn}
\bibitem{You:2023rmn}
Z.~Q.~You, Z.~Yi and Y.~Wu,
%``Constraints on primordial curvature power spectrum with pulsar timing arrays,''
[arXiv:2307.04419 [gr-qc]].
%0 citations counted in INSPIRE as of 14 Jul 2023


%\cite{Wang:2023div}
\bibitem{Wang:2023div}
S.~Wang and Z.~C.~Zhao,
%``Unveiling the Graviton Mass Bounds through Analysis of 2023 Pulsar Timing Array Datasets,''
[arXiv:2307.04680 [astro-ph.HE]].
%0 citations counted in INSPIRE as of 15 Jul 2023



%\cite{Figueroa:2023zhu}
\bibitem{Figueroa:2023zhu}
D.~G.~Figueroa, M.~Pieroni, A.~Ricciardone and P.~Simakachorn,
%``Cosmological Background Interpretation of Pulsar Timing Array Data,''
[arXiv:2307.02399 [astro-ph.CO]].
%14 citations counted in INSPIRE as of 20 Jul 2023


%\cite{Choudhury:2023kam}
\bibitem{Choudhury:2023kam}
S.~Choudhury,
%``Single field inflation in the light of NANOGrav 15-year Data: Quintessential interpretation of blue tilted tensor spectrum through Non-Bunch Davies initial condition,''
[arXiv:2307.03249 [astro-ph.CO]].
%3 citations counted in INSPIRE as of 20 Jul 2023


%\cite{HosseiniMansoori:2023mqh}
\bibitem{HosseiniMansoori:2023mqh}
S.~A.~Hosseini Mansoori, F.~Felegray, A.~Talebian and M.~Sami,
%``PBHs and GWs from $\mathbb{T}^2$-inflation and NANOGrav 15-year data,''
[arXiv:2307.06757 [astro-ph.CO]].
%1 citations counted in INSPIRE as of 20 Jul 2023


%\cite{Ge:2023rce}
\bibitem{Ge:2023rce}
S.~Ge,
%``Stochastic gravitational wave background: birth from axionic string-wall death,''
[arXiv:2307.08185 [gr-qc]].
%1 citations counted in INSPIRE as of 30 Jul 2023

%\cite{Bian:2023dnv}
\bibitem{Bian:2023dnv}
L.~Bian, S.~Ge, J.~Shu, B.~Wang, X.~Y.~Yang and J.~Zong,
%``Gravitational wave sources for Pulsar Timing Arrays,''
[arXiv:2307.02376 [astro-ph.HE]].
%10 citations counted in INSPIRE as of 30 Jul 2023

%\cite{Kawasaki:2023rfx}
\bibitem{Kawasaki:2023rfx}
M.~Kawasaki and K.~Murai,
%``Enhancement of gravitational waves at Q-ball decay including non-linear density perturbations,''
[arXiv:2308.13134 [astro-ph.CO]].
%1 citations counted in INSPIRE as of 29 Aug 2023





%\cite{Yi:2023tdk}
\bibitem{Yi:2023tdk}
Z.~Yi, Z.~Q.~You and Y.~Wu,
%``Model-independent reconstruction of the primordial curvature power spectrum from PTA data,''
[arXiv:2308.05632 [astro-ph.CO]].
%3 citations counted in INSPIRE as of 29 Aug 2023


%\cite{An:2023jxf}
\bibitem{An:2023jxf}
H.~An, B.~Su, H.~Tai, L.~T.~Wang and C.~Yang,
%``Phase transition during inflation and the gravitational wave signal at pulsar timing arrays,''
[arXiv:2308.00070 [astro-ph.CO]].
%6 citations counted in INSPIRE as of 29 Aug 2023


%\cite{Zhang:2023nrs}
\bibitem{Zhang:2023nrs}
Z.~Zhang, C.~Cai, Y.~H.~Su, S.~Wang, Z.~H.~Yu and H.~H.~Zhang,
%``Nano-Hertz gravitational waves from collapsing domain walls associated with freeze-in dark matter in light of pulsar timing array observations,''
[arXiv:2307.11495 [hep-ph]].
%14 citations counted in INSPIRE as of 29 Aug 2023

%\cite{DiBari:2023upq}
\bibitem{DiBari:2023upq}
P.~Di Bari and M.~H.~Rahat,
%``The split majoron model confronts the NANOGrav signal,''
[arXiv:2307.03184 [hep-ph]].
%14 citations counted in INSPIRE as of 29 Aug 2023


%\cite{Jiang:2023qbm}
\bibitem{Jiang:2023qbm}
S.~Jiang, A.~Yang, J.~Ma and F.~P.~Huang,
%``Implication of nano-Hertz stochastic gravitational wave on dynamical dark matter through a first-order phase transition,''
[arXiv:2306.17827 [hep-ph]].
%17 citations counted in INSPIRE as of 29 Aug 2023


%\cite{Bhattacharya:2023ysp}
\bibitem{Bhattacharya:2023ysp}
G.~Bhattacharya, S.~Choudhury, K.~Dey, S.~Ghosh, A.~Karde and
N.~S.~Mishra,
%``Evading no-go for PBH formation and production of SIGWs using Multiple Sharp Transitions in EFT of single field inflation,''
[arXiv:2309.00973 [astro-ph.CO]].
%0 citations counted in INSPIRE as of 06 Sep 2023




%\cite{Choudhury:2023hfm}
\bibitem{Choudhury:2023hfm}
S.~Choudhury, A.~Karde, S.~Panda and M.~Sami,
%``Scalar induced gravity waves from ultra slow-roll Galileon inflation,''
[arXiv:2308.09273 [astro-ph.CO]].
%3 citations counted in INSPIRE as of 06 Sep 2023



%\cite{Bringmann:2023opz}
\bibitem{Bringmann:2023opz}
T.~Bringmann, P.~F.~Depta, T.~Konstandin, K.~Schmidt-Hoberg and
C.~Tasillo,
%``Does NANOGrav observe a dark sector phase transition?,''
[arXiv:2306.09411 [astro-ph.CO]].
%0 citations counted in INSPIRE as of 22 Jun 2023

%\cite{Choudhury:2023hvf}
\bibitem{Choudhury:2023hvf}
S.~Choudhury, S.~Panda and M.~Sami,
%``Galileon inflation evades the no-go for PBH formation in the single-field framework,''
JCAP \textbf{08} (2023), 078 doi:10.1088/1475-7516/2023/08/078
[arXiv:2304.04065 [astro-ph.CO]].
%23 citations counted in INSPIRE as of 12 Sep 2023


%\cite{Choudhury:2023kdb}
\bibitem{Choudhury:2023kdb}
S.~Choudhury, A.~Karde, S.~Panda and M.~Sami,
%``Primordial non-Gaussianity from ultra slow-roll Galileon inflation,''
[arXiv:2306.12334 [astro-ph.CO]].
%9 citations counted in INSPIRE as of 12 Sep 2023


%\cite{Huang:2023chx}
\bibitem{Huang:2023chx}
H.~L.~Huang, Y.~Cai, J.~Q.~Jiang, J.~Zhang and Y.~S.~Piao,
%``Supermassive primordial black holes in multiverse: for nano-Hertz gravitational wave and high-redshift JWST galaxies,''
[arXiv:2306.17577 [gr-qc]].
%20 citations counted in INSPIRE as of 12 Sep 2023


%\cite{Jiang:2023gfe}
\bibitem{Jiang:2023gfe}
J.~Q.~Jiang, Y.~Cai, G.~Ye and Y.~S.~Piao,
%``Broken blue-tilted inflationary gravitational waves: a joint analysis of NANOGrav 15-year and BICEP/Keck 2018 data,''
[arXiv:2307.15547 [astro-ph.CO]].
%8 citations counted in INSPIRE as of 12 Sep 2023

%\cite{Zhu:2023lbf}
\bibitem{Zhu:2023lbf}
M.~Zhu, G.~Ye and Y.~Cai,
%``Pulsar timing array observations as a possible hint for nonsingular cosmology,''
[arXiv:2307.16211 [astro-ph.CO]].
%3 citations counted in INSPIRE as of 12 Sep 2023

%\cite{Ben-Dayan:2023lwd}
\bibitem{Ben-Dayan:2023lwd}
I.~Ben-Dayan, U.~Kumar, U.~Thattarampilly and A.~Verma,
%``Probing The Early Universe Cosmology With NANOGrav: Possibilities and Limitations,''
[arXiv:2307.15123 [astro-ph.CO]].
%7 citations counted in INSPIRE as of 12 Sep 2023




%\cite{Franciolini:2023pbf}
\bibitem{Franciolini:2023pbf}
G.~Franciolini, A.~Iovino, Junior., V.~Vaskonen and H.~Veermae,
%``The recent gravitational wave observation by pulsar timing arrays and primordial black holes: the importance of non-gaussianities,''
[arXiv:2306.17149 [astro-ph.CO]].
%64 citations counted in INSPIRE as of 12 Sep 2023


%\cite{Ellis:2023oxs}
\bibitem{Ellis:2023oxs}
J.~Ellis, M.~Fairbairn, G.~Franciolini, G.~H\"utsi, A.~Iovino,
M.~Lewicki, M.~Raidal, J.~Urrutia, V.~Vaskonen and H.~Veerm\"ae,
%``What is the source of the PTA GW signal?,''
[arXiv:2308.08546 [astro-ph.CO]].
%9 citations counted in INSPIRE as of 12 Sep 2023



%\cite{Liu:2023ymk}
\bibitem{Liu:2023ymk}
L.~Liu, Z.~C.~Chen and Q.~G.~Huang,
%``Implications for the non-Gaussianity of curvature perturbation from pulsar timing arrays,''
[arXiv:2307.01102 [astro-ph.CO]].
%39 citations counted in INSPIRE as of 14 Sep 2023

%\cite{Liu:2023pau}
\bibitem{Liu:2023pau}
L.~Liu, Z.~C.~Chen and Q.~G.~Huang,
%``Probing the equation of state of the early Universe with pulsar timing arrays,''
[arXiv:2307.14911 [astro-ph.CO]].
%14 citations counted in INSPIRE as of 14 Sep 2023


%\cite{Madge:2023cak}
\bibitem{Madge:2023cak}
E.~Madge, E.~Morgante, C.~Puchades-Ib\'a\~nez, N.~Ramberg,
W.~Ratzinger, S.~Schenk and P.~Schwaller,
%``Primordial gravitational waves in the nano-Hertz regime and PTA data -- towards solving the GW inverse problem,''
[arXiv:2306.14856 [hep-ph]].
%40 citations counted in INSPIRE as of 14 Sep 2023


%\cite{Huang:2023zvs}
\bibitem{Huang:2023zvs}
M.~X.~Huang, F.~Wang and Y.~K.~Zhang,
%``The Interplay Between the Muon $g-2$ Anomaly and the PTA nHZ Gravitational Waves from Domain Walls in NMSSM,''
[arXiv:2309.06378 [hep-ph]].
%0 citations counted in INSPIRE as of 14 Sep 2023


%\cite{Fu:2023aab}
\bibitem{Fu:2023aab}
C.~Fu, J.~Liu, X.~Y.~Yang, W.~W.~Yu and Y.~Zhang,
%``Explaining Pulsar Timing Array Observations with Primordial Gravitational Waves in Parity-Violating Gravity,''
[arXiv:2308.15329 [astro-ph.CO]].
%2 citations counted in INSPIRE as of 14 Sep 2023


%\cite{Maji:2023fhv}
\bibitem{Maji:2023fhv}
R.~Maji and W.~I.~Park,
%``Supersymmetric $U(1)_{B-L}$ flat direction and NANOGrav 15 year data,''
[arXiv:2308.11439 [hep-ph]].
%3 citations counted in INSPIRE as of 14 Sep 2023

%\cite{Gangopadhyay:2023qjr}
\bibitem{Gangopadhyay:2023qjr}
M.~R.~Gangopadhyay, V.~V.~Godithi, K.~Ichiki, R.~Inui, T.~Kajino,
A.~Manusankar, G.~J.~Mathews and Yogesh,
%``Is NanoGRAV signals pointing towards resonant particle creation during inflation?,''
[arXiv:2309.03101 [astro-ph.CO]].
%1 citations counted in INSPIRE as of 16 Sep 2023




%\cite{Schwaller:2015tja}
\bibitem{Schwaller:2015tja}
P.~Schwaller,
%``Gravitational Waves from a Dark Phase Transition,''
Phys. Rev. Lett. \textbf{115} (2015) no.18, 181101
doi:10.1103/PhysRevLett.115.181101 [arXiv:1504.07263 [hep-ph]].
%240 citations counted in INSPIRE as of 04 Jul 2023




%\cite{Ratzinger:2020koh}
\bibitem{Ratzinger:2020koh}
W.~Ratzinger and P.~Schwaller,
%``Whispers from the dark side: Confronting light new physics with NANOGrav data,''
SciPost Phys. \textbf{10} (2021) no.2, 047
doi:10.21468/SciPostPhys.10.2.047 [arXiv:2009.11875
[astro-ph.CO]].
%86 citations counted in INSPIRE as of 04 Jul 2023



%\cite{Ashoorioon:2022raz}
\bibitem{Ashoorioon:2022raz}
A.~Ashoorioon, K.~Rezazadeh and A.~Rostami,
%``NANOGrav signal from the end of inflation and the LIGO mass and heavier primordial black holes,''
Phys. Lett. B \textbf{835} (2022), 137542
doi:10.1016/j.physletb.2022.137542 [arXiv:2202.01131
[astro-ph.CO]].
%39 citations counted in INSPIRE as of 04 Jul 2023

%\cite{Choudhury:2023vuj}
\bibitem{Choudhury:2023vuj}
S.~Choudhury, M.~R.~Gangopadhyay and M.~Sami,
%``No-go for the formation of heavy mass Primordial Black Holes in Single Field Inflation,''
[arXiv:2301.10000 [astro-ph.CO]].
%36 citations counted in INSPIRE as of 20 Jul 2023

%\cite{Choudhury:2023jlt}
\bibitem{Choudhury:2023jlt}
S.~Choudhury, S.~Panda and M.~Sami,
%``No-go for PBH formation in EFT of single field inflation,''
[arXiv:2302.05655 [astro-ph.CO]].
%32 citations counted in INSPIRE as of 20 Jul 2023


%\cite{Choudhury:2023rks}
\bibitem{Choudhury:2023rks}
S.~Choudhury, S.~Panda and M.~Sami,
%``Quantum loop effects on the power spectrum and constraints on primordial black holes,''
[arXiv:2303.06066 [astro-ph.CO]].
%29 citations counted in INSPIRE as of 20 Jul 2023


%\cite{Bian:2022qbh}
\bibitem{Bian:2022qbh}
L.~Bian, S.~Ge, C.~Li, J.~Shu and J.~Zong,
%``Domain Wall Network: A Dual Solution for Gravitational Waves and Hubble Tension?,''
[arXiv:2212.07871 [hep-ph]].
%9 citations counted in INSPIRE as of 30 Jul 2023






%\cite{Machado:2018nqk}
\bibitem{Machado:2018nqk}
C.~S.~Machado, W.~Ratzinger, P.~Schwaller and B.~A.~Stefanek,
%``Audible Axions,''
JHEP \textbf{01} (2019), 053 doi:10.1007/JHEP01(2019)053
[arXiv:1811.01950 [hep-ph]].
%54 citations counted in INSPIRE as of 04 Jul 2023

%\cite{Vagnozzi:2020gtf}
\bibitem{Vagnozzi:2020gtf}
S.~Vagnozzi,
%``Implications of the NANOGrav results for inflation,''
Mon. Not. Roy. Astron. Soc. \textbf{502} (2021) no.1, L11-L15
doi:10.1093/mnrasl/slaa203 [arXiv:2009.13432 [astro-ph.CO]].
%156 citations counted in INSPIRE as of 20 Jan 2024




%\cite{Kuroyanagi:2020sfw}
\bibitem{Kuroyanagi:2020sfw}
S.~Kuroyanagi, T.~Takahashi and S.~Yokoyama,
%``Blue-tilted inflationary tensor spectrum and reheating in the light of NANOGrav results,''
JCAP \textbf{01} (2021), 071 doi:10.1088/1475-7516/2021/01/071
[arXiv:2011.03323 [astro-ph.CO]].
%55 citations counted in INSPIRE as of 20 Jan 2024



%\cite{Benetti:2021uea}
\bibitem{Benetti:2021uea}
M.~Benetti, L.~L.~Graef and S.~Vagnozzi,
%``Primordial gravitational waves from NANOGrav: A broken power-law approach,''
Phys. Rev. D \textbf{105} (2022) no.4, 043520
doi:10.1103/PhysRevD.105.043520 [arXiv:2111.04758 [astro-ph.CO]].
%88 citations counted in INSPIRE as of 20 Jan 2024







%\cite{Kamali:2020drm}
\bibitem{Kamali:2020drm}
V.~Kamali and R.~Brandenberger,
%``Creating spatial flatness by combining string gas cosmology and power law inflation,''
Phys. Rev. D \textbf{101} (2020) no.10, 103512
doi:10.1103/PhysRevD.101.103512 [arXiv:2002.09771 [hep-th]].
%3 citations counted in INSPIRE as of 20 Jul 2023



%\cite{Brandenberger:2015kga}
\bibitem{Brandenberger:2015kga}
R.~H.~Brandenberger,
%``String Gas Cosmology after Planck,''
Class. Quant. Grav. \textbf{32} (2015) no.23, 234002
doi:10.1088/0264-9381/32/23/234002 [arXiv:1505.02381 [hep-th]].
%34 citations counted in INSPIRE as of 20 Jul 2023



%\cite{Brandenberger:2006pr}
\bibitem{Brandenberger:2006pr}
R.~H.~Brandenberger, S.~Kanno, J.~Soda, D.~A.~Easson, J.~Khoury,
P.~Martineau, A.~Nayeri and S.~P.~Patil,
%``More on the spectrum of perturbations in string gas cosmology,''
JCAP \textbf{11} (2006), 009 doi:10.1088/1475-7516/2006/11/009
[arXiv:hep-th/0608186 [hep-th]].
%70 citations counted in INSPIRE as of 20 Jul 2023


%\cite{Ashtekar:2011ni}
\bibitem{Ashtekar:2011ni}
A.~Ashtekar and P.~Singh,
%``Loop Quantum Cosmology: A Status Report,''
Class. Quant. Grav. \textbf{28} (2011), 213001
doi:10.1088/0264-9381/28/21/213001 [arXiv:1108.0893 [gr-qc]].
%927 citations counted in INSPIRE as of 06 Sep 2023





%\cite{Bojowald:2011iq}
\bibitem{Bojowald:2011iq}
M.~Bojowald, G.~Calcagni and S.~Tsujikawa,
%``Observational test of inflation in loop quantum cosmology,''
JCAP \textbf{11} (2011), 046 doi:10.1088/1475-7516/2011/11/046
[arXiv:1107.1540 [gr-qc]].
%86 citations counted in INSPIRE as of 06 Sep 2023


%\cite{Mielczarek:2009vi}
\bibitem{Mielczarek:2009vi}
J.~Mielczarek,
%``Tensor power spectrum with holonomy corrections in LQC,''
Phys. Rev. D \textbf{79} (2009), 123520
doi:10.1103/PhysRevD.79.123520 [arXiv:0902.2490 [gr-qc]].
%34 citations counted in INSPIRE as of 06 Sep 2023


%\cite{Bojowald:2008ik}
\bibitem{Bojowald:2008ik}
M.~Bojowald,
%``Consistent Loop Quantum Cosmology,''
Class. Quant. Grav. \textbf{26} (2009), 075020
doi:10.1088/0264-9381/26/7/075020 [arXiv:0811.4129 [gr-qc]].
%103 citations counted in INSPIRE as of 06 Sep 2023







%\cite{Calcagni:2020tvw}
\bibitem{Calcagni:2020tvw}
G.~Calcagni and S.~Kuroyanagi,
%``Stochastic gravitational-wave background in quantum gravity,''
JCAP \textbf{03} (2021), 019 doi:10.1088/1475-7516/2021/03/019
[arXiv:2012.00170 [gr-qc]].
%29 citations counted in INSPIRE as of 20 Jul 2023



%\cite{Koshelev:2020foq}
\bibitem{Koshelev:2020foq}
A.~S.~Koshelev, K.~Sravan Kumar, A.~Mazumdar and
A.~A.~Starobinsky,
%``Non-Gaussianities and tensor-to-scalar ratio in non-local R$^{2}$-like inflation,''
JHEP \textbf{06} (2020), 152 doi:10.1007/JHEP06(2020)152
[arXiv:2003.00629 [hep-th]].
%38 citations counted in INSPIRE as of 20 Jul 2023


%\cite{Koshelev:2017tvv}
\bibitem{Koshelev:2017tvv}
A.~S.~Koshelev, K.~Sravan Kumar and A.~A.~Starobinsky,
%``$R^2$ inflation to probe non-perturbative quantum gravity,''
JHEP \textbf{03} (2018), 071 doi:10.1007/JHEP03(2018)071
[arXiv:1711.08864 [hep-th]].
%73 citations counted in INSPIRE as of 20 Jul 2023

%\cite{Baumgart:2021ptt}
\bibitem{Baumgart:2021ptt}
M.~Baumgart, J.~J.~Heckman and L.~Thomas,
%``CFTs blueshift tensor fluctuations universally,''
JCAP \textbf{07} (2022) no.07, 034
doi:10.1088/1475-7516/2022/07/034 [arXiv:2109.08166 [hep-ph]].
%6 citations counted in INSPIRE as of 15 Aug 2023



%\cite{Hwang:2005hb}
\bibitem{Hwang:2005hb}
  J.~c.~Hwang and H.~Noh,
  %``Classical evolution and quantum generation in generalized gravity theories including string corrections and tachyon: Unified analyses,''
  Phys.\ Rev.\ D {\bf 71} (2005) 063536
  doi:10.1103/PhysRevD.71.063536
  [gr-qc/0412126].
  %%CITATION = doi:10.1103/PhysRevD.71.063536;%%
  %87 citations counted in INSPIRE as of 11 Jun 2018


%\cite{Nojiri:2006je}
\bibitem{Nojiri:2006je}
  S.~Nojiri, S.~D.~Odintsov and M.~Sami,
  %``Dark energy cosmology from higher-order, string-inspired gravity and its reconstruction,''
  Phys.\ Rev.\ D {\bf 74} (2006) 046004
  doi:10.1103/PhysRevD.74.046004
  [hep-th/0605039].
  %%CITATION = doi:10.1103/PhysRevD.74.046004;%%
  %271 citations counted in INSPIRE as of 15 Jul 2019




%\cite{Cognola:2006sp}
\bibitem{Cognola:2006sp}
  G.~Cognola, E.~Elizalde, S.~Nojiri, S.~Odintsov and S.~Zerbini,
  %``String-inspired Gauss-Bonnet gravity reconstructed from the universe expansion history and yielding the transition from matter dominance to dark energy,''
  Phys.\ Rev.\ D {\bf 75} (2007) 086002
  doi:10.1103/PhysRevD.75.086002
  [hep-th/0611198].
  %%CITATION = doi:10.1103/PhysRevD.75.086002;%%
  %163 citations counted in INSPIRE as of 12 Jun 2018



%\cite{Nojiri:2005vv}
\bibitem{Nojiri:2005vv}
  S.~Nojiri, S.~D.~Odintsov and M.~Sasaki,
  %``Gauss-Bonnet dark energy,''
  Phys.\ Rev.\ D {\bf 71} (2005) 123509
  doi:10.1103/PhysRevD.71.123509
  [hep-th/0504052].
  %%CITATION = doi:10.1103/PhysRevD.71.123509;%%
  %435 citations counted in INSPIRE as of 12 Jun 2018


%\cite{Nojiri:2005jg}
\bibitem{Nojiri:2005jg}
  S.~Nojiri and S.~D.~Odintsov,
  %``Modified Gauss-Bonnet theory as gravitational alternative for dark energy,''
  Phys.\ Lett.\ B {\bf 631} (2005) 1
  doi:10.1016/j.physletb.2005.10.010
  [hep-th/0508049].
  %%CITATION = doi:10.1016/j.physletb.2005.10.010;%%
  %478 citations counted in INSPIRE as of 12 Jun 2018







  %\cite{Satoh:2007gn}
\bibitem{Satoh:2007gn}
  M.~Satoh, S.~Kanno and J.~Soda,
  %``Circular Polarization of Primordial Gravitational Waves in String-inspired Inflationary Cosmology,''
  Phys.\ Rev.\ D {\bf 77} (2008) 023526
  doi:10.1103/PhysRevD.77.023526
  [arXiv:0706.3585 [astro-ph]].
  %%CITATION = doi:10.1103/PhysRevD.77.023526;%%
  %64 citations counted in INSPIRE as of 16 Aug 2018



%\cite{Kanti:2015pda}
\bibitem{Kanti:2015pda}
 P.~Kanti, R.~Gannouji and N.~Dadhich,
 %``Gauss-Bonnet Inflation,''
 Phys.\ Rev.\ D {\bf 92} (2015) no.4, 041302
 doi:10.1103/PhysRevD.92.041302
 [arXiv:1503.01579 [hep-th]].
 %%CITATION = doi:10.1103/PhysRevD.92.041302;%%
 %27 citations counted in INSPIRE as of 12 Jun 2018


%\cite{Kanti:1998jd}
\bibitem{Kanti:1998jd}
 P.~Kanti, J.~Rizos and K.~Tamvakis,
 %``Singularity free cosmological solutions in quadratic gravity,''
 Phys.\ Rev.\ D {\bf 59} (1999) 083512
 doi:10.1103/PhysRevD.59.083512
 [gr-qc/9806085].
 %%CITATION = doi:10.1103/PhysRevD.59.083512;%%
 %79 citations counted in INSPIRE as of 12 Jun 2018

%\cite{Kleihaus:2019rbg}
\bibitem{Kleihaus:2019rbg}
B.~Kleihaus, J.~Kunz and P.~Kanti,
%``Particle-like ultracompact objects in Einstein-scalar-Gauss-Bonnet theories,''
Phys. Lett. B \textbf{804} (2020), 135401
doi:10.1016/j.physletb.2020.135401 [arXiv:1910.02121 [gr-qc]].
%25 citations counted in INSPIRE as of 28 Jul 2021




 %\cite{Bakopoulos:2019tvc}
\bibitem{Bakopoulos:2019tvc}
 A.~Bakopoulos, P.~Kanti and N.~Pappas,
 %``Existence of solutions with a horizon in pure scalar-Gauss-Bonnet theories,''
 Phys.\ Rev.\ D {\bf 101} (2020) no.4, 044026
 doi:10.1103/PhysRevD.101.044026
 [arXiv:1910.14637 [hep-th]].
 %%CITATION = doi:10.1103/PhysRevD.101.044026;%%
 %2 citations counted in INSPIRE as of 25 Mar 2020


%\cite{Odintsov:2020sqy}
\bibitem{Odintsov:2020sqy}
S.~D.~Odintsov, V.~K.~Oikonomou and F.~P.~Fronimos,
%``Rectifying Einstein-Gauss-Bonnet Inflation in View of GW170817,''
Nucl. Phys. B \textbf{958} (2020), 115135
doi:10.1016/j.nuclphysb.2020.115135 [arXiv:2003.13724 [gr-qc]].
%70 citations counted in INSPIRE as of 07 Jan 2022


%\cite{Oikonomou:2021kql}
\bibitem{Oikonomou:2021kql}
V.~K.~Oikonomou,
%``A refined Einstein\textendash{}Gauss\textendash{}Bonnet inflationary theoretical framework,''
Class. Quant. Grav. \textbf{38} (2021) no.19, 195025
doi:10.1088/1361-6382/ac2168 [arXiv:2108.10460 [gr-qc]].
%2 citations counted in INSPIRE as of 07 Jan 2022



\bibitem{reviews1}
 S.~Nojiri, S.~D.~Odintsov and V.~K.~Oikonomou,
  %``Modified Gravity Theories on a Nutshell: Inflation, Bounce and Late-time Evolution,''
  Phys.\ Rept.\  {\bf 692} (2017) 1
  %doi:10.1016/j.physrep.2017.06.001
  [arXiv:1705.11098 [gr-qc]].
  %%CITATION = %doi:10.1016/j.physrep.2017.06.001;%%
  %21 citations counted in INSPIRE as of 20 Aug 2017

\bibitem{reviews2}


 S. Capozziello, M. De Laurentis,
   %``Extended Theories of Gravity,''
   Phys.\ Rept.\  {\bf 509}, 167 (2011);\\
   %[arXiv:1108.6266 [gr-qc]].
   %%CITATION = ARXIV:1108.6266;%%
 V.~Faraoni and S.~Capozziello,
  %``Beyond Einstein Gravity : A Survey of Gravitational Theories for Cosmology and Astrophysics,''
  Fundam.\ Theor.\ Phys.\  {\bf 170} (2010).
  %%doi:10.1007/978-94-007-0165-6
  %%CITATION = %doi:10.1007/978-94-007-0165-6;%%
  %64 citations counted in INSPIRE as of 16 Sep 2017



\bibitem{reviews3}
S. Nojiri, S.D. Odintsov,
  %``Introduction to modified gravity and gravitational alternative for dark
  %energy,''
  eConf {\bf C0602061}, 06 (2006)
  [Int.\ J.\ Geom.\ Meth.\ Mod.\ Phys.\  {\bf 4}, 115 (2007)].
  %[arXiv:hep-th/0601213];
  %%CITATION = 00436,4,115;%%


   \bibitem{reviews4}

S. Nojiri, S.D. Odintsov,
   %``Unified cosmic history in modified gravity: from F(R) theory to
   %Lorentz non-invariant models,''
   Phys.\ Rept.\  {\bf 505}, 59 (2011);
   %[arXiv:1011.0544 [gr-qc]].
   %%CITATION = ARXIV:1011.0544;%%

%\cite{NANOGrav:2023hvm}
\bibitem{NANOGrav:2023hvm}
A.~Afzal \textit{et al.} [NANOGrav],
%``The NANOGrav 15 yr Data Set: Search for Signals from New Physics,''
Astrophys. J. Lett. \textbf{951} (2023) no.1, L11
doi:10.3847/2041-8213/acdc91 [arXiv:2306.16219 [astro-ph.HE]].
%282 citations counted in INSPIRE as of 26 Jan 2024



\end{thebibliography}

\documentclass[aps,prd,onecolumn,groupedaddress,showpacs,nofootinbib,amssymb]{revtex4}
\usepackage[dvips]{graphicx}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{graphicx,color}
\usepackage{amsfonts}
\usepackage{bm}
\usepackage{cancel}
\usepackage{comment}

\newcommand\be{\begin{equation}}
\newcommand\ee{\end{equation}}

\allowdisplaybreaks[4]

\begin{document}

\title{Red or Blue Tensor Spectrum from GW170817-compatible Einstein-Gauss-Bonnet Theory: A Detailed Analysis}
\author{V.K. Oikonomou,$^{1,2}$}\,
\email{v.k.oikonomou1979@gmail.com;voikonomou@gapps.auth.gr}
\author{Pyotr Tsyba$^{2}$}\,
\email{pyotrtsyba@gmail.com}
\author{Olga Razina$^{2}$}\,
\email{olvikraz@mail.ru} \affiliation{$^{1)}$ Department of
Physics, Aristotle University of Thessaloniki, Thessaloniki 54124,
Greece\\ $^{2)}$L.N. Gumilyov Eurasian National University -
Astana, 010008, Kazakhstan}

\tolerance=5000

\begin{abstract}
In this work we shall prove that the tensor spectral index of the
primordial tensor perturbations for GW170817-compatible
Einstein-Gauss-Bonnet theories, takes the approximate simplified
form $n_{\mathcal{T}}\simeq 2\left(-1+\frac{1}{\lambda(\phi)}
\right)\epsilon_1$ at leading order, with $\lambda (\phi)$ being a
function of the scalar field which depends on the scalar field
potential and the second derivative of the scalar-Gauss-Bonnet
coupling $\xi''(\phi)$. With our analysis we aim to provide a
definitive criterion for selecting Einstein-Gauss-Bonnet models
that can provide a blue-tilted inflationary phenomenology, by
simply looking at the scalar potential and the scalar-Gauss-Bonnet
coupling. We shall prove this using two distinct approaches and as
we show the tilt of the tensor spectral index is determined by the
values of the potential $V(\phi)$ and of scalar-Gauss-Bonnet
coupling at first horizon crossing. Specifically the blue-tilted
tensor spectral index can occur when $\xi''(\phi_*)V(\phi_*)>0$ at
first horizon crossing.
\end{abstract}

\pacs{04.50.Kd, 95.36.+x, 98.80.-k, 98.80.Cq,11.25.-w}

\maketitle

\section{Introduction}

The current interest in theoretical physics is on inflationary
theories
\cite{inflation1,inflation2,inflation3,inflation4,inflation5}
which will be experimentally scrutinized in the next decade.
Indeed, several future experiments will directly seek for hints of
the inflationary theory, either direct like the stage 4 Cosmic
Microwave Background (CMB) experiments
\cite{CMB-S4:2016ple,SimonsObservatory:2019qwx} which will seek
for $B$-modes of inflation, or indirect like the future
gravitational waves experiments
\cite{Hild:2010id,Baker:2019nia,Smith:2019wny,Crowder:2005nr,Smith:2016jqs,Seto:2001qf,Kawamura:2020pcg,Bull:2018lat,LISACosmologyWorkingGroup:2022jok}.
These experiments, apart from constraining, it is also possible to
rule out inflation \cite{Vagnozzi:2022qmc}. The gravitational wave
experiments will be able to discover the existence or not of a
stochastic gravitational waves background with small anisotropies,
which cannot be attributed to astrophysical sources due to the
small anisotropies.

In 2023 an exciting observation from the NANOGrav collaboration
\cite{nanograv} has placed many hopes on the stochastic
gravitational wave background existence. Specifically NANOGrav
\cite{nanograv} confirmed the existence of a stochastic
gravitational wave background and this observation was also
confirmed simultaneously by other pulsar timing arrays (PTA)
experiments \cite{Antoniadis:2023ott,Reardon:2023gzh,Xu:2023wog}.
The absence of large anisotropies \cite{NANOGrav:2023gor}, the
absence of a single supermassive black hole merger event and the
lack of a definitive theoretical solution to the last parsec
problem \cite{Sampson:2015ada} makes the cosmological
interpretation of the signal more likely, at lest for the time
being and many works have appeared in the literature, explaining
the 2023 NANOGrav signal
\cite{sunnynew,Oikonomou:2023qfz,Cai:2023dls,Han:2023olf,Guo:2023hyp,Yang:2023aak,Addazi:2023jvg,Li:2023bxy,Niu:2023bsr,Yang:2023qlf,Datta:2023vbs,Du:2023qvj,Salvio:2023ynn,Yi:2023mbm,You:2023rmn,Wang:2023div,Figueroa:2023zhu,Choudhury:2023kam,HosseiniMansoori:2023mqh,Ge:2023rce,Bian:2023dnv,Kawasaki:2023rfx,Yi:2023tdk,An:2023jxf,Zhang:2023nrs,DiBari:2023upq,Jiang:2023qbm,Bhattacharya:2023ysp,Choudhury:2023hfm,Bringmann:2023opz,Choudhury:2023hvf,Choudhury:2023kdb,Huang:2023chx,Jiang:2023gfe,Zhu:2023lbf,Ben-Dayan:2023lwd,Franciolini:2023pbf,Ellis:2023oxs,Liu:2023ymk,Liu:2023pau,Madge:2023cak,Huang:2023zvs,Fu:2023aab,Maji:2023fhv,Gangopadhyay:2023qjr},
see also
\cite{Schwaller:2015tja,Ratzinger:2020koh,Ashoorioon:2022raz,Choudhury:2023vuj,Choudhury:2023jlt,Choudhury:2023rks,Bian:2022qbh}
for previous works, and also
\cite{Guo:2023hyp,Yang:2023aak,Machado:2018nqk}. In the case that
the stochastic gravitational wave background is due to a
cosmological source, the inflationary perspective of generating
such a signal is significantly constrained because in such a case
a significantly blue-tilted tensor spectral index combined with a
low-reheating temperature is required
\cite{sunnynew,Oikonomou:2023qfz,Vagnozzi:2020gtf,Kuroyanagi:2020sfw,Benetti:2021uea}.

In fact, in some cases it is not necessary to have a strongly
blue-tilted tensor spectral index to explain the NANOGrav signal,
but an abnormal reheating or a post-reheating era is required
\cite{Oikonomou:2023qfz}. The blue-tilted inflationary tensor
spectral index however is a prerequisite to explain the NANOGrav
theories, and there exist theories that can generate a blue-tilted
tensor spectral index, like for example string gas theories
\cite{Kamali:2020drm,Brandenberger:2015kga,Brandenberger:2006pr},
Loop Quantum Cosmology scenarios
\cite{Ashtekar:2011ni,Bojowald:2011iq,Mielczarek:2009vi,Bojowald:2008ik},
non-local gravity models
\cite{Calcagni:2020tvw,Koshelev:2020foq,Koshelev:2017tvv} or
conformal field theories \cite{Baumgart:2021ptt}. One
string-originating theory that can produce a blue-tilted tensor
spectral index is Einstein-Gauss-Bonnet theory, and for mainstream
of articles and reviews see Refs.
\cite{Hwang:2005hb,Nojiri:2006je,Cognola:2006sp,Nojiri:2005vv,Nojiri:2005jg,Satoh:2007gn,Kanti:2015pda,Kanti:1998jd,Kleihaus:2019rbg,Odintsov:2020sqy,Oikonomou:2021kql}
and \cite{reviews1,reviews2,reviews3,reviews4,inflation5}
respectively.

Due to the importance of obtaining a blue-tilted spectral index,
in this article we shall derive and prove in detail that the
tensor spectral index for a general GW170817-compatible
Einstein-Gauss-Bonnet theory can be written in a closed and
simplified form for two distinct approaches, and we shall analyze
the cases when it is blue-tilted or red-tilted and which
parameters or functions determine the sign of the tensor spectral
index at first horizon crossing. As we will show, the size and
magnitude of the second derivative of the Gauss-Bonnet scalar
coupling function $\xi''(\phi)$ at first horizon crossing and the
value of the scalar field potential $V(\phi)$, fully determine the
sign of the tensor spectral index.

\subsection{Inflation with GW170817-compatible Einstein-Gauss-Bonnet Theory and the Tensor Spectral Index}

Before we go to the core of our analysis, let us discuss the
inflationary dynamics of GW170817-compatible Einstein-Gauss-Bonnet
gravity which was developed in
\cite{Odintsov:2020sqy,Oikonomou:2021kql}. The
Einstein-Gauss-Bonnet gravity action is the following,
\begin{equation}
\label{action} \centering
S=\int{d^4x\sqrt{-g}\left(\frac{R}{2\kappa^2}-\frac{\omega}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{1}{2}\xi(\phi)\mathcal{G}\right)}\,
,
\end{equation}
where $R$ stands for the Ricci scalar, also $\kappa=\frac{1}{M_p}$
where $M_p$ is the reduced Planck mass and furthermore
$\mathcal{G}$ stands for the four dimensional Gauss-Bonnet
invariant which is
$\mathcal{G}=R^2-4R_{\alpha\beta}R^{\alpha\beta}+R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$
where $R_{\alpha\beta}$ is the Ricci and also
$R_{\alpha\beta\gamma\delta}$ is the Riemann tensor. Let us
consider a flat Friedmann-Robertson-Walker (FRW) geometric
background, with metric,
\begin{equation}
\label{metric} \centering
ds^2=-dt^2+a(t)^2\sum_{i=1}^{3}{(dx^{i})^2}\, ,
\end{equation}
with $a(t)$ denoting as usual the scale factor. By varying the
gravitational action with respect to the metric and with respect
to a time-dependent scalar field, for a FRW metric, the field
equations read,
\begin{equation}
\label{motion1} \centering
\frac{3H^2}{\kappa^2}=\frac{1}{2}\dot\phi^2+V+12 \dot\xi H^3\, ,
\end{equation}
\begin{equation}
\label{motion2} \centering \frac{2\dot
H}{\kappa^2}=-\dot\phi^2+4\ddot\xi H^2+8\dot\xi H\dot H-4\dot\xi
H^3\, ,
\end{equation}
\begin{equation}
\label{motion3} \centering \ddot\phi+3H\dot\phi+V'+12 \xi'H^2(\dot
H+H^2)=0\, .
\end{equation}
For the inflationary era analysis, we shall assume that the
following slow-roll conditions apply,
\begin{equation}\label{slowrollhubble}
\dot{H}\ll H^2,\,\,\ \frac{\dot\phi^2}{2} \ll V,\,\,\,\ddot\phi\ll
3 H\dot\phi\, .
\end{equation}
The speed of the tensor perturbations of the FRW background metric
is \cite{Hwang:2005hb,Odintsov:2020sqy,Oikonomou:2021kql},
\begin{equation}
\label{GW} \centering c_T^2=1-\frac{Q_f}{2Q_t}\, ,
\end{equation}
with $Q_f=8 (\ddot\xi-H\dot\xi)$, $Q_t=F+\frac{Q_b}{2}$,
$F=\frac{1}{\kappa^2}$ and $Q_b=-8 \dot\xi H$. The GW170817 event
imposed the constraint that the gravitational wave speed is nearly
equal to the speed of light, so this means that $Q_f\simeq 0$
which in turn means that $\ddot\xi\simeq H\dot\xi$. By expressing
this in terms of the scalar field, we have,
\begin{equation}
\label{constraint1} \centering
\xi''\dot\phi^2+\xi'\ddot\phi=H\xi'\dot\phi\, ,
\end{equation}
with the ``prime'' indicating differentiation with respect to the
scalar field. Furthermore, if we require that the following
condition holds true,
\begin{equation}\label{firstslowroll}
 \xi'\ddot\phi \ll\xi''\dot\phi^2\, ,
\end{equation}
we can rewrite the constraint (\ref{constraint1}) in the following
way,
\begin{equation}
\label{constraint} \centering
\dot{\phi}\simeq\frac{H\xi'}{\xi''}\, ,
\end{equation}
therefore by combining Eqs. (\ref{motion3}) and (\ref{constraint})
we obtain,
\begin{equation}
\label{motion4} \centering \frac{\xi'}{\xi''}\simeq-\frac{1}{3
H^2}\left(V'+12 \xi'H^4\right)\, .
\end{equation}
Now we further assume that
\begin{equation}\label{mainnewassumption}
\kappa \frac{\xi '}{\xi''}\ll 1\, ,
\end{equation}
\begin{equation}\label{scalarfieldslowrollextra}
12 \dot\xi H^3=12 \frac{\xi'^2H^4}{\xi''}\ll V\, ,
\end{equation}
which will simplify the field equations and we can obtain the
slow-roll indices in closed form. Hence, in view of Eqs.
(\ref{slowrollhubble}), (\ref{constraint}) and
(\ref{scalarfieldslowrollextra}), the field equations become,
\begin{equation}
\label{motion5} \centering H^2\simeq\frac{\kappa^2V}{3}\, ,
\end{equation}
\begin{equation}
\label{motion6} \centering \dot H\simeq-\frac{1}{2}\kappa^2
\dot\phi^2\, ,
\end{equation}
\begin{equation}
\label{motion8} \centering \dot\phi\simeq\frac{H\xi'}{\xi''}\, ,
\end{equation}
and furthermore the condition (\ref{scalarfieldslowrollextra})
becomes,
\begin{equation}\label{mainconstraint2}
 \frac{4\kappa^4\xi'^2V}{3\xi''}\ll 1\, .
\end{equation}
In addition, the differential equation (\ref{motion4}) which
directly relates the functional forms of the  scalar field
potential and the scalar field coupling function of the
Gauss-Bonnet invariant, takes the form,
\begin{equation}
\label{maindiffeqnnew} \centering
\frac{V'}{V^2}+\frac{4\kappa^4}{3}\xi'\simeq 0\, .
\end{equation}
In view of the above, we can obtain the slow-roll indices acquire
a closed and simplified form,
\begin{equation}
\label{index1} \centering \epsilon_1\simeq\frac{\kappa^2\omega
}{2}\left(\frac{\xi'}{\xi''}\right)^2\,
,\,\,\,\epsilon_2\simeq1-\epsilon_1-\frac{\xi'\xi'''}{\xi''^2},\,\,\,
\epsilon_3=0\, ,
\end{equation}
\begin{equation}
\label{index4} \centering
\epsilon_4\simeq\frac{\xi'}{2\xi''}\frac{\mathcal{E}'}{\mathcal{E}}\,
,
\end{equation}
\begin{equation}
\label{index5} \centering
\epsilon_5\simeq-\frac{2\kappa^4\xi'^2V}{3\xi''-4\kappa^4\xi'^2V}\,
,
\end{equation}
\begin{equation}
\label{index6} \centering \epsilon_6\simeq
\epsilon_5(1-\epsilon_1)\, ,
\end{equation}
with $\mathcal{E}=\mathcal{E}(\phi)$ being defined in the
following way,
\begin{equation}\label{functionE}
\mathcal{E}(\phi)=\frac{1}{\kappa^2}\left(
1+72\frac{\epsilon_1^2}{\lambda^2} \right)\, ,
\end{equation}
with the function $\lambda(\phi)$ being defined as follows,
\begin{equation}\label{lambdadef}
\lambda(\phi)=\frac{3\omega}{4\xi''\kappa^2 V}\, .
\end{equation}
Now, in terms of the slow-roll indices, the scalar spectral index
$n_{\mathcal{S}}$, the tensor-to-scalar ratio $r$ and the
tensor-spectral index $n_{\mathcal{T}}$ are,
\begin{equation}
\label{spectralindex} \centering
n_{\mathcal{S}}=1-4\epsilon_1-2\epsilon_2-2\epsilon_4\, ,
\end{equation}
\begin{equation}\label{tensorspectralindex}
n_{\mathcal{T}}=-2\left( \epsilon_1+\epsilon_6 \right)\, ,
\end{equation}
\begin{equation}\label{tensortoscalar}
r=16\left|\left(\frac{\kappa^2Q_e}{4H}-\epsilon_1\right)\frac{2c_A^3}{2+\kappa^2Q_b}\right|\,
,
\end{equation}
where $c_A$ denotes the sound wave speed,
\begin{equation}
\label{sound} \centering c_A^2=1+\frac{Q_aQ_e}{3Q_a^2+
\dot\phi^2(\frac{2}{\kappa^2}+Q_b)}\, ,
\end{equation}
with,
\begin{align}\label{qis}
& Q_a=-4 \dot\xi H^2,\,\,\,Q_b=-8 \dot\xi H,\,\,\,
Q_t=F+\frac{Q_b}{2},\,\,\, Q_c=0,\,\,\,Q_e=-16 \dot{\xi}\dot{H}\,
.
\end{align}
The tensor-to-scalar ratio can be further simplified as follows,
\begin{equation}\label{tensortoscalarratiofinal}
r\simeq 16\epsilon_1\, .
\end{equation}
Regarding the tensor spectral index, we shall demonstrate that it
can be written in the following closed and simple form,
\begin{equation}\label{tensorspectralindexfinal}
n_{\mathcal{T}}\simeq -2\epsilon_1\left ( 1-\frac{1}{\lambda
}+\frac{\epsilon_1}{\lambda}\right)\, ,
\end{equation}
where $\lambda(\phi)$ is defined in Eq. (\ref{lambdadef}). We
shall show that the above form can be obtained by using various
approaches. To start with, let us begin our analysis from the
tensor spectral index of Eq. (\ref{tensorspectralindex}) and focus
on the slow-roll index $\epsilon_6$ defined in Eq. (\ref{index6}).
So we start from,
\begin{equation}\label{epsilon6index1}
\epsilon_6=-\frac{2\kappa^4\xi'^2\left(1-\frac{1}{2}\kappa^2\omega
\left(\frac{\xi'}{\xi''} \right)^2
\right)}{\xi''^2\left(\frac{3}{\xi''V}-4\kappa^4\left(\frac{\xi'}{\xi''}\right)^2
\right)}\, ,
\end{equation}
which can be written in terms of the slow-roll index $\epsilon_1$
as follows,
\begin{equation}\label{epsilon6index2}
\epsilon_6=-\frac{\frac{4\kappa^2}{\omega}\epsilon_1\left(1-\epsilon_1
\right)}{\frac{3}{\xi''V}-\frac{8\kappa^2\epsilon_1}{\omega}}\, ,
\end{equation}
or equivalently,
\begin{equation}\label{epsilon6index2}
\epsilon_6=-\frac{\epsilon_1(1-\epsilon_1)}{\frac{3\omega}{4\kappa^2\xi''V}-2\epsilon_1}
\, .
\end{equation}
Thus by inserting the above in Eq. (\ref{tensorspectralindex}) we
get,
\begin{equation}\label{tensorspectralindexfinal1}
n_{\mathcal{T}}\simeq -2\left
(\epsilon_1-\frac{\epsilon_1(1-\epsilon_1)}{\frac{3\omega}{4\kappa^2\xi''}-2\epsilon_1}
\right)\, ,
\end{equation}
and since $\epsilon_1\ll 1$ we expand the above for small values
of $\epsilon_1$ and we get,
\begin{equation}\label{tensorspectralindexfinal2}
n_{\mathcal{T}}\simeq 2\left(-1+\frac{1}{\lambda}
\right)\epsilon_1-\frac{2(\lambda-2)}{\lambda^2}\epsilon_1^2+...-2^{n-1}\frac{\lambda-2}{\lambda^n}\epsilon_1^n\,
.
\end{equation}
Thus at first order, the tensor spectral index has the following
form,
\begin{equation}\label{firstordertensorspectralindex}
n_{\mathcal{T}}\simeq 2\left(-1+\frac{1}{\lambda}
\right)\epsilon_1\, .
\end{equation}
We can arrive at the same expression by using a different
approach, starting from Eq. (\ref{tensorspectralindex}) and using
Eq. (\ref{index6}), and then we get approximately,
\begin{equation}\label{oneneweqn}
n_{\mathcal{T}}\simeq
-2\left(\epsilon_1+\epsilon_5-\epsilon_5\epsilon_1 \right)\, ,
\end{equation}
and it can be proven that $\epsilon_5$ takes a simpler form than
the one in Eq. (\ref{index5}), since,
\begin{equation}\label{rewrittemepsilon5}
\epsilon_5=-\frac{\epsilon_1}{\frac{3\omega}{\xi''V4\kappa^2}-2\epsilon_1}\,
.
\end{equation}
Due to the approximation we assumed earlier, namely
$\frac{4\xi'^2\kappa^4V}{3\xi''}\ll 1$, we get,
\begin{equation}\label{approxepsilon5}
\epsilon_5\simeq
-\frac{2\kappa^4\xi'^2V}{3\xi''}=-\frac{\epsilon_1}{\lambda}\, ,
\end{equation}
thus, the tensor spectral index (\ref{oneneweqn}) acquires the
following form,
\begin{equation}\label{tensoapproach2final}
n_{\mathcal{T}}\simeq
-2\epsilon_1(1-\frac{1}{\lambda})-\frac{2\epsilon_1^2}{\lambda}\,
,
\end{equation}
which is identical to the expression given in Eq.
(\ref{firstordertensorspectralindex}) if terms of the order $\sim
\mathcal{O}(\epsilon_1)$ are kept in the expansion. Thus we proven
using two distinct approaches that the tensor spectral index at
first order approximation can take the simplified form given in
Eq. (\ref{firstordertensorspectralindex}). It is apparent that the
tilt of the spectral index is critically affected by the values of
the function $\lambda(\phi)$ at first horizon crossing, since in
our formalism $\epsilon_1$ is always positive. Thus the following
cases determine the tilt of the tensor spectral index,
\begin{equation}\label{interactionterm2}
n_{\mathcal{T}}\simeq\left\{
\begin{array}{c}
  -2\epsilon_1,\,\,\,\mathrm{when}\,\,\,|\lambda(\phi_*)|\gg 1  \\
 \frac{2\epsilon_1}{\lambda},\,\,\,\mathrm{when}\,\,\,|\lambda(\phi_*)|\ll
 1\\
  0,\,\,\,\mathrm{when}\,\,\,|\lambda(\phi_*)|\simeq 1 \, ,\\
  \frac{\epsilon_1(1-\lambda)}{\lambda}>0,\,\,\,\mathrm{when}\,\,\,\lambda(\phi_*)<1 \, ,
\end{array}\right.
\end{equation}
where $\phi_*$ is the value of the scalar field at the first
horizon crossing. Thus the strongly blue-tilted tensor spectral
index can occur when $|\lambda(\phi_*)|\ll 1$ and when $\lambda
(\phi_*)>0$, and also in the case that $\lambda<1$ the tensor
spectral index is also blue-tilted but not necessarily strongly
blue-tilted. Now it is apparent that the values and the sign of
the function $\lambda (\phi)$ are determined by the values and
signs of $\xi''(\phi)$ and of $V(\phi)$ at the first horizon
crossing. Thus the blue-tilted tensor spectral index can occur
when $\xi''(\phi_*)V(\phi_*)>0$ at first horizon crossing. Clearly
this is more or less model dependent, but our analysis provides
insights in the analysis of inflationary phenomenology of
Einstein-Gauss-Bonnet models, because models with positive
potentials must have $\xi''(\phi)>0$ in order to generate a
blue-tilted tensor spectral index and also models with negative
values of the potential at first horizon crossing must have
$\xi''(\phi_*)<0$ in order to generate a blue-tilted tensor
spectral index.

Let us consider in brief some illustrative examples, the
phenomenology of which is well studied in the literature and these
provide a viable inflationary phenomenology
\cite{Oikonomou:2021kql}. Consider the following functional form
of the scalar coupling function $\xi(\phi)$,
\begin{equation}
\label{modelApower} \xi(\phi)=\beta  \left(\frac{\phi
}{M}\right)^{\nu }\, ,
\end{equation}
with $\beta$ being a dimensionless parameter, and $M$ is a free
parameter with mass dimensions $[m]^1$. By combining Eqs.
(\ref{modelApower}) and (\ref{maindiffeqnnew}) and solving the
differential equation (\ref{maindiffeqnnew}), we get the following
scalar field potential,
\begin{equation}
\label{potApower} \centering V(\phi)=\frac{3}{3 \gamma  \kappa
^4+4 \beta  \kappa ^4 \left(\frac{\phi }{M}\right)^{\nu }} \, ,
\end{equation}
with $\gamma$ being a dimensionless integration constant. Since in
this case, we always have $\xi''(\phi)>0$, the sign of the tensor
spectral index will be determined by the scalar potential value,
which is positive in order for a viable phenomenology to be
ensured. Indeed a viable phenomenology is guaranteed for the
following values of the free parameters \cite{Oikonomou:2021kql}
$\mu=2.09\times 10^{-22}\times \kappa$,
$\beta=[0.000025,1.69055]$, $\gamma=10^{200}$, for $N=[50,63]$
$e$-foldings, for which the corresponding tensor spectral index
values are in the range $n_{\mathcal{T}}=[0.006,1.108]$. Thus for
these values the scalar potential is positive, and as we expected,
the tensor spectral index is blue-tilted. Another model of this
sort is obtained for the scalar coupling function $\xi(\phi)$
having the form $\xi(\phi)=\beta  \exp \left(\left(\frac{\phi
}{M}\right)^2\right)$, where $\beta$ is a dimensionless parameter,
and $M$ is a free parameter with mass dimensions $[m]^1$. By
combining $\xi (\phi)$ and Eq. (\ref{maindiffeqnnew}) and by
solving the differential equation (\ref{maindiffeqnnew}) we obtain
the following scalar potential,
\begin{equation}
\label{potA} \centering V(\phi)=\frac{3}{3 \gamma  \kappa ^4+4
\beta  \kappa ^4 e^{\frac{\phi ^2}{M^2}}} \, ,
\end{equation}
where $\gamma$ is a dimensionless integration constant. In this
case we have $\xi''(\phi)>0$ if $\beta>0$ and $\xi''(\phi)<0$ if
$\beta<0$, and the same applies for the scalar potential, which is
however affected by the free parameter $\gamma$ too. Thus in this
case, the product $\xi(\phi_*)V(\phi_*)$ at the first horizon
crossing will determine the sign of the tensor spectral index.
Therefore, one must find which values of the free parameters
guarantee a viable inflationary phenomenology, and the condition
$\xi(\phi_*)V(\phi_*)>0$ can easily be examined. In this case, a
viable phenomenology is guaranteed for the following values of the
free parameters \cite{Oikonomou:2021kql} $\mu=[22.091,22.0914]$,
$\beta=-1.5$, $\gamma=2$, for $N=60$ $e$-foldings. Thus both the
potential and the second derivative of the scalar Gauss-Bonnet
coupling function are negative, thus the tensor spectral index is
positive. Indeed, as it was shown in \cite{Oikonomou:2021kql} for
these values of the free parameters we have
$n_{\mathcal{T}}=[0.378,0.379]$, hence it is positive and also one
may verify that $\xi(\phi_*)V(\phi_*)>0$. Therefore, with our
analysis one may know beforehand which Einstein-Gauss-Bonnet model
may yield a blue-tilted tensor spectral index. However, we need to
note that while the examples we presented can predict a blue
tensor tilt, it is not clear whether those, or other model in the
class, can provide the high values of the tensor spectral index
required to explain PTA data \cite{sunnynew,NANOGrav:2023hvm}.
This feature can be model dependent, so it is left as a question
to the reader and for future work.

\section{Conclusions}

In this paper we focused on the issue when the tensor spectral
index of the primordial tensor perturbations for
GW170817-compatible Einstein-Gauss-Bonnet theories is blue-tilted
or red-tilted. With our analysis we aimed to provide a definitive
 and simple criterion for appropriately selecting Einstein-Gauss-Bonnet models that can
provide a blue-tilted inflationary phenomenology, by simply
selecting appropriately the scalar potential and the
scalar-Gauss-Bonnet coupling. The importance of a blue-tilted
inflationary phenomenology is great, after the 2023 NANOGrav
observation of the stochastic gravitational wave background. Using
the GW170817-compatible formalism of Einstein-Gauss-Bonnet
inflationary theories, we showed, by using two distinct
approaches, that at leading order, the tensor spectral index takes
the approximate simplified form $n_{\mathcal{T}}\simeq
2\left(-1+\frac{1}{\lambda(\phi)} \right)\epsilon_1$ at leading
order, with $\lambda (\phi)$ being a function of the scalar field
which depends on the scalar field potential and the second
derivative of the scalar-Gauss-Bonnet coupling $\xi''(\phi)$.
Since in our formalism, the slow-roll index $\epsilon_1$ is always
positive, the tilt of the spectral index depends on the sign of
the function $\lambda (\phi)$ at the first horizon crossing, which
in turn depends on the values of the potential $V(\phi)$ and of
scalar-Gauss-Bonnet coupling at first horizon crossing. As we
showed, when $\lambda (\phi_*)\ll 1$ and when $\lambda (\phi_*)<
1$, the tensor spectral index is blue-tilted, while when
$\lambda(\phi_*)\simeq 1$, and $\lambda(\phi_*)> 1$ the tilt is
red, where $\phi_*$ the value of the scalar field at the first
horizon crossing. In terms of the scalar field potential and the
scalar-Gauss-Bonnet coupling, the blue-tilted tensor spectral
index can occur when $\xi''(\phi_*)V(\phi_*)>0$ at first horizon
crossing. Hence our analysis provides insights in the study of
Einstein-Gauss-Bonnet inflationary theories, making it possible to
chose beforehand blue-tilted inflationary phenomenologies.

\section*{Acknowledgments}

This research has been is funded by the Committee of Science of
the Ministry of Education and Science of the Republic of
Kazakhstan (Grant No. AP14869238).

\end{document}